OpenStudy (anonymous):

Solve the following linear equation. Describe the solution set using interval notation.

OpenStudy (anonymous):

\[\frac{u-6}{7} \ge \frac{2u-1}{3}\]

OpenStudy (anonymous):

\[(-\infty,\frac{1}{11}]\]

OpenStudy (anonymous):

Can you show me how you came about solving this?

OpenStudy (anonymous):

made a mistake..

OpenStudy (anonymous):